OpenStudy (slipknot306):

Can someone please help me with math??? I will give a medal!!!!!! -8[(12-6)+3]+4 --------------- 9-2(2+3) It is like a fraction

OpenStudy (alexandervonhumboldt2):

-68/35

OpenStudy (anonymous):

-8[(12-6)+3]+4=-8[6+3]+4=-8(9)+4=-72+4=-68

OpenStudy (anonymous):

9-2(2+3)=9-2(5)=9-10=-1

OpenStudy (anonymous):

so answer is -68/-1=68

OpenStudy (slipknot306):

Thank you so much

OpenStudy (slipknot306):

What about 1+[2x-3(1.8/0.2)]

OpenStudy (anonymous):

what is x?? value of x is given??

OpenStudy (anonymous):

1.8/0.2=18/2=9 so 3(1.8/0.2)=3*9=27

OpenStudy (slipknot306):

it does not say

OpenStudy (anonymous):

if the value of x is not given, required solution =1+2x-27=2x-26

OpenStudy (slipknot306):

Thank you sooooo much